# Stochastic Multi-round Submodular Optimization with Budget

### v1

@@ 4.5 Full Proof of Theorem 1
We conclude the proof by using all the above inequalities as follows:
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-\sigma({\sf MultiGr}_{\delta,\xi}(I))$ (28)
$=\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right]$
$=\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\overline{\bm{\Delta}}_{t,i}\right]$
$\ \ \ \ \ \ \ +\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\overline{\bm{\Delta}}_{t,i}\right]-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\overline{\bm{\Delta}}_{t,i}\right]+\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\overline{\bm{\Delta}}_{t,i}\right]-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\tilde{\bm{\Delta}}_{t,i}\right]$
$\ \ \ \ \ \ \ +\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\tilde{\bm{\Delta}}_{t,i}\right]-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right]$
$=\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}(\Delta_{t,i}-\overline{\bm{\Delta}}_{t,i})\right]$
@@
$\ \ \ \ \ \ \ +\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}(\tilde{\bm{\Delta}}_{t,i}-\Delta_{t,i})\right]$
$\leq\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}(\Delta_{t,i}-\overline{\bm{\Delta}}_{t,i})\right]+\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}(\tilde{\bm{\Delta}}_{t,i}-\Delta_{t,i})\right]$
[…]
$\leq B(2\delta+\xi\Lambda)+B\xi\Lambda+B\delta+B\xi\Lambda$
$=3B(\delta+\xi\Lambda).$
By (28), the claim follows. ∎
@@
By definition of $\lambda$, there exists $t\in[T]$ and $v\in V$ such that $\lambda\leq\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]$. By the optimality of policy $\pi^{*}$, we necessarily have $\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]\leq\sigma(\pi^{*})$. Thus, by combining the obtained inequalities, the claim of the lemma follows. ∎
By combining Lemma 1 with (8), we get $\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))\geq\sigma(\pi^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-2B(\delta+\Lambda\xi)$, that is
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))\geq\frac{\sigma(\pi^{*})}{2}-B(\delta+\Lambda\xi).$ (29)
Then, we get
$\sigma({\sf MultiGr}_{\delta,\xi}(I))$ $\geq\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-3B(\delta+\xi\Lambda)$ (30)
$\geq\frac{1}{2}\sigma(\pi^{*})-B(\delta+\Lambda\xi)-3B(\delta+\xi\Lambda)$ (31)
$=\frac{1}{2}\sigma(\pi^{*})-4\delta(1+\Lambda)B$
$\geq\frac{1}{2}\sigma(\pi^{*})-\frac{\delta(1+\Lambda)B}{c}$
$\geq\frac{1}{2}\sigma(\pi^{*})-\epsilon\cdot\lambda$ (32)
$\geq\frac{1}{2}\sigma(\pi^{*})-\epsilon\cdot\sigma(\pi^{*})$ (33)
@@
$=\left(\frac{1}{2}-\epsilon\right)\sigma(\pi^{*}),$ (34)
where (30) follows from Lemma 2, (31) follows from (29), (32) holds since we previously set $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$ with $c\leq 1/4$, and (33) holds by Lemma 3. By (34), the theorem follows.
@@ 5 Conclusions
Our analysis aims to provide theoretical guarantees on the approximation that can be achieved in polynomial time. We proposed two algorithms: one based on a dynamic programming approach, that is able to compute the optimal multi-round adaptive policy in polynomial time whenever the single-round optimization problem can be efficiently solved, and a simpler greedy algorithm that, despite being only semi-adaptive, still returns a constant approximation of the optimum. Even if the proposed algorithms are essentially simple (mainly the second one that follows a greedy programming technique) and they are proved to run in polynomial time, they involve some routines, such as Monte Carlo simulations, that are known to limit their practical application. Anyway, we believe that, as occurred for many similar problems, our algorithms may foster further research towards more practical approaches, that, e.g., replace the more expensive routines with heuristics. Not only, but our algorithm will also serve as a benchmark for the evaluation of these heuristics.
Apart for optimizing the algorithm for practical application, a theoretically relevant open question is that of providing better approximation algorithms for the multi-round optimization problem studied in this work. We point out that, for submodular single-round optimization problems under budget constraints, there is a $(1-1/e-\epsilon)$ approximation algorithm [22], and cannot be substantially improved in light of the $(1-1/e+\epsilon)$-inapproximability shown by [19]. Thus, it would be interesting to close the gap between $(1-1/e-\epsilon)$ and $(1-1/2-\epsilon)$ on the approximation to the multi-round optimization problem.
Finally, we recall that our algorithms assume that state distributions are known in advance, even if their realization is unknown. It would be undoubtedly interesting to explore the existence of algorithms with good approximation guarantees even if the state distributions must be learnt or they are provided by a some learning framework (without any precision guarantee).

### v2

@@ 4.5 Full Proof of Theorem 1
We recall that $A$ denotes the event “$|\Delta_{t,i}-\tilde{\bm{\Delta}}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$” and that $\mathbb{P}[A]\geq 1-\xi$; furthermore, let $\overline{A}$ denote the complementary set of $A$, so that $\mathbb{P}[\overline{A}]\leq\xi$. We conclude the proof by using all the above inequalities as follows:
@@
By (27), the claim follows. ∎ By exploiting the polynomial boundedness property, the following lemma provides a lower bound on the optimal adaptive value.
@@
Proof of Lemma 3.
By definition of $\lambda$, there exist $t\in[T]$ and $v\in V$ such that $\lambda\leq\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]$. By the optimality of policy $\pi^{*}$, we necessarily have $\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]\leq\sigma(\pi^{*})$. Thus, by combining the obtained inequalities, the claim of the lemma follows. ∎
Finally, putting together the above results, we obtain
$\sigma({\sf MultiGr}_{\delta,\xi}(I))\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-3B(\delta+\xi\Lambda)\geq\left(1-\frac{1}{e}\right)\sigma(\pi^{*})-4B(\delta+\xi\Lambda)$
$\ \geq\left(1-\frac{1}{e}\right)\sigma(\pi^{*})-\left(\frac{4B(\delta+\xi\Lambda)}{\lambda}\right)\sigma(\pi^{*})\geq\left(1-\frac{1}{e}\right)\sigma(\pi^{*})-\left(\frac{B(\delta+\xi\Lambda)}{c\lambda}\right)\sigma(\pi^{*})=\left(1-\frac{1}{e}-\epsilon\right)\sigma(\pi^{*}),$
where the first inequality follows from Lemma 2, the second one holds by (19), the third one follows from Lemma 3, the fourth one holds since $40<c\leq 1/$ 4, the last equality holds since $\delta=\xi=\frac{\lambda c\epsilon}{B(1+\Lambda)}$ and shows the desired approximation factor.
5 Budget-adaptivity Gap
In this section, we introduce and quantify the budget-adaptivity gap, to measure how much a fully adaptive optimal policy is better than an optimal partially adaptive policy that, as in the greedy multi-round policy considered above (see Section 4.4), allocates the budget in advance.
Definition 3.
Given an instance $I$ of SBMSm, the budget-adaptivity gap of $I$ is defined as $BGap(I):=\sigma(\pi^{*})/\sigma(\pi)$, where $\pi^{*}$ is an optimal adaptive policy for $I$, and $\pi$ is a policy that is optimal among all partially adaptive policies in which we first choose the budget $b_{t}$ to be spent for each round $t$ (i.e., non-adaptively) and then we adaptively select $b_{t}$ items at each round $t$. Furthermore, the budget-adaptivity gap of SBMSm is defined as $BGap:=\sup_{I\in\text{SBMSm}}BGap(I)$.
In the following theorem we provide a tight constant bound on the adpativity gap, thus showing that the fully adaptive optimum is, in general, strictly higher than the partially adaptive one, but it can still be approximated well by resorting to partial adaptivity.
Theorem 2.
The budget-adaptivity gap of SBMSm is $e/(e-1)\approx 1.582$.
@@ 5 Budget-adaptivity Gap
We first show that $BGap\leq e/(e-1)$. Let $I$ be an arbitrary input instance of SBMSm, $\pi^{*}$ be an optimal adaptive policy for $I$ and $\pi$ be an optimal partially adaptive for $I$ (that allocates the budget in advance); furthermore, let ${\sf MultiGr}_{0,0}(I)$ denote the multi-round greedy policy defined in Section 4.4, assuming that the probabilistic oracles (Oracle1 and Oracle2 of Section 4.1) works with $\delta=\xi=0$, that is, they are able to return with probability $1$ the exact expected increment of the objective functions. If $\delta=\xi=0$, the same analysis as that provided in Theorem 1 holds with $\epsilon=0$, thus showing that the approximation factor of ${\sf MultiGr}_{0,0}(I)$ is $1-1/e=(e-1)/e$. Thus, we have $\frac{1}{BGap(I)}=\frac{\sigma(\pi)}{\sigma(\pi^{*})}\geq\frac{\sigma({\sf MultiGr}_{0,0}(I))}{\sigma(\pi^{*})}\geq\frac{e-1}{e},$ that is, $BGap(I)\leq e/(e-1)$, and by the arbitrariness of $I$ we obtain $BGap\leq e/(e-1)$.
To show the claim, it remains to prove that $BGap\geq e/(e-1)$. We equivalently show that, for any $\epsilon>0$, there exists an instance $I(\epsilon)$ of SBMSm parametrized by $\epsilon$ such that $BGap(I(\epsilon))>e/(e-1)-\epsilon$; then, by the arbitrariness of $\epsilon$ we obtain $BGap\geq e/(e-1)$.
For a fixed $\epsilon>0$, consider an instance $I(\epsilon)$ of the multi-round binary stochastic probing problem (see Section 2.5) defined as follows: we have $T$ rounds and a set $V$ of $n:=T^{3/2}$ items, where $T:=T(\epsilon)$ is a sufficiently large perfect square that will be fixed later; the budget is $B:=n=T^{3/2}$; each item $v\in V$ is active with probability $p:=1/\sqrt{T}$, for each round $t\in[T]$; the objective function $g_{t}$ of each round $t$ is equal to $1$ if the number of active items is at least $1$, and 00 otherwise.
In the following, we will provide a lower bound on $BGap(I(\epsilon))$, by comparing the value of an optimal partially adaptive policy $\pi$ and a specific fully adaptive policy $\pi^{\prime}$ defined below. If we show that $\sigma(\pi^{\prime})/\sigma(\pi)>\frac{e}{e-1}-\epsilon$, this implies that $BGap(I(\epsilon))\geq\sigma(\pi^{\prime})/\sigma(\pi)>\frac{e}{e-1}-\epsilon$, that is, the desired lower bound.
Let $\pi$ be the partially adaptive policy that non-adaptively selects the first $B/T=\sqrt{T}$ items at each round. Let $\pi^{\prime}$ be the fully adaptive policy that, at each round, repedeatly select all items until finding an active one, and then move to the subsequent round; the execution ends if all the budget has been spent or if there are no more rounds available. We have that
$\sigma(\pi)$ $=\sum_{t\in[T]}\mathbb{P}[\text{$\pi$ has selected at least an active item at round $t$}]$
$=\sum_{t\in[T]}\left(1-\mathbb{P}[\text{$\pi$ has selected no active item at round $t$}]\right)$
$=\sum_{t\in[T]}\left(1-\prod_{k=1}^{B/T}\mathbb{P}[\text{the $k$-th item selected at round $t$ by $\pi$ is not active}]\right)$ (28)
$=\sum_{t\in[T]}\left(1-\prod_{k=1}^{\sqrt{T}}(1-p)\right)=T\cdot\left(1-\left(1-\frac{1}{\sqrt{T}}\right)^{\sqrt{T}}\right),$ (29)
where (28) holds by the independence of the items activation process.
Now, we determine a lower bound on $\sigma(\pi^{\prime})$. For any $k\in[B]$, let $\bm{X}_{k}$ be an independent Bernoulli variable with parameter $p$. We observe that $\min\{T,\sum_{k\in[B]}\bm{X}_{k}\}$ is distributed as the value of policy $\pi^{\prime}$. Indeed, $\pi^{\prime}$ is equivalent, in terms of total revenue, to a single-round policy that repeatedly selects $B$ items, each one independently active with probability $p=1/\sqrt{T}$, counts the number of selected active items and returns the maximum between such number and $T$.
For $t:=\sqrt{T}$, we have
$\sigma(\pi^{\prime})$ $\geq\mathbb{P}[\text{the value of $\pi^{\prime}$ is at least $T-t$}]\cdot(T-t)=\mathbb{P}[\text{$\pi^{\prime}$ selects at least $T-t$ active items}]\cdot(T-t)$
$=\mathbb{P}\left[\min\left\{T,\sum_{k\in[B]}\bm{X}_{k}\right\}\geq T-t\right]\cdot(T-t)$ (30)
$=\mathbb{P}\left[\sum_{k\in[B]}\bm{X}_{k}\geq T-t\right]\cdot(T-t)=\mathbb{P}\left[\sum_{k\in[B]}\bm{X}_{k}-\mathbb{E}\left[\sum_{k\in[B]}\bm{X}_{k}\right]\geq-t\right]\cdot(T-t)$ (31)
$\geq\left(1-\frac{1}{e^{2t^{2}/B}}\right)\cdot(T-t)$ (32)
$=\left(1-\frac{1}{e^{2/\sqrt{T}}}\right)\cdot(T-\sqrt{T}),$ (33)
where (30) holds by the above observations, (31) follows from $\mathbb{E}\left[\sum_{k\in[B]}\bm{X}_{k}\right]=Bp=T^{3/2}\cdot T^{-1/2}=T$, (32) holds by the Hoeffding’s inequality [31] and (33) follows from $t=\sqrt{T}$ and $B=T^{3/2}$. By combining (29) and (33) we obtain
$\frac{\sigma(\pi^{\prime})}{\sigma(\pi)}\geq\frac{\left(1-\frac{1}{e^{2/\sqrt{T}}}\right)\cdot(T-\sqrt{T})}{\left(1-\left(1-\frac{1}{\sqrt{T}}\right)^{\sqrt{T}}\right)\cdot T}.$ (34)
As the right-hand part of (34) tends to $\frac{1}{1-\frac{1}{e}}=\frac{e}{e-1}$ for $T$ tending to $\infty$, by taking a sufficiently large perfect square $T:=T(\epsilon)$ in instance $I(\epsilon)$ we obtain
@@
$\frac{\sigma(\pi^{\prime})}{\sigma(\pi)}\geq\frac{\left(1-\frac{1}{e^{2/\sqrt{T}}}\right)\cdot(T-\sqrt{T})}{\left(1-\left(1-\frac{1}{\sqrt{T}}\right)^{\sqrt{T}}\right)\cdot T}>\frac{e}{e-1}-\epsilon.$
By the initial observations, the above inequality implies that the budget adaptivity gap of $I(\epsilon)$ is higher than $\frac{e}{e-1}-\epsilon$, and by the arbitrariness of $\epsilon$ the claim follows. ∎
[…]
Corollary 4.
The approximation factor of the greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}$ is at least $1-\frac{1}{e}$, for any $\delta,\xi>0$.
Proof.
Assume by contradiction that the approximation factor $\alpha$ of ${\sf MultiGr}_{\delta,\xi}$ is higher than $1-\frac{1}{e}$, for some $\delta,\xi>0$. Let $\epsilon:=\frac{e}{e-1}-\frac{1}{\alpha}>0$ be a fixed constant, $I(\epsilon)$ be the input instance considered in the proof of Theorem 2, $\pi^{*}$ and $\pi$ be respectively an optimal fully adaptive policy and an optimal partially adaptive policy for $I(\epsilon)$. In Theorem 2 we showed that the budget adaptivity gap of $I(\epsilon)$ is strictly higher than $\frac{e}{e-1}-\epsilon=\frac{1}{\alpha}$. Then, as ${\sf MultiGr}_{\delta,\xi}$ is a partially adaptive greedy policy, we necessarily have that $\sigma(\pi)\geq\sigma({\sf MultiGr}_{\delta,\xi}(I(\epsilon)))$. We conclude that $\alpha>\frac{1}{BGap(I(\epsilon))}=\frac{\sigma(\pi)}{\sigma(\pi^{*})}\geq\frac{\sigma({\sf MultiGr}_{\delta,\xi}(I(\epsilon)))}{\sigma(\pi^{*})},$ but this contradicts the $\alpha$-approximation assumed at the beginning.
@@ 6 Conclusions
Our analysis aims to provide theoretical guarantees on the approximation that can be achieved in polynomial time. We proposed two algorithms: one based on a dynamic programming approach, that is able to compute the optimal multi-round adaptive policy in polynomial time whenever a single-round optimization problem associated with each round can be efficiently solved (e.g., if the number of items is constant and the number of random realizations is polynomial), and a simpler greedy algorithm that, despite being only partially adaptive, still returns a $(1-1/e-\epsilon)$-approximation of the optimum, and such result is essentially optimal in light of the hardness result of [23] (holding for simpler non-adaptive single-round maximization problems). Furthermore, by introducing the budget-adaptivity gap, we measured how much is lost, in terms of value obtained, if we distribute the budget over all rounds in advance, i.e. in a non-adaptive way.
A first research direction is that of combining multiple rounds, budget and adaptivity in other classical optimization problems, as we did in submodular maximization problems, and then study computational complexity, approximation and budget-adaptivity gap.
Another interesting open question is related to the efficiency of the above greedy algorithm in terms of run-time. Indeed, even if the proposed approximation algorithm we provided is essentially simple and it is proved to run in polynomial time, it involve some routines, such as Monte Carlo simulations, that are known to limit their practical application. Anyway, we believe that, as occurred for many similar problems, our algorithm may foster further research towards more practical approaches, that, e.g., replace the more expensive routines with heuristics. Not only, but our algorithm will also serve as a benchmark for the evaluation of these heuristics.
Finally, we recall that the algorithms designed in this work assume that state distributions are known in advance, even if their realization is unknown. It would be undoubtedly interesting to explore the existence of algorithms with good approximation guarantees even if the state distributions must be learnt or they are provided by a some learning framework (without any precision guarantee).